Thursday, August 18, 2016

Collimator triplet oil change!

This morning the guys drained the collimator triplet & replaced the lens fluid in the 2 gaps.  Like last time, the scene was quite gruesome!  The problematic original lens fluid (which has vastly different specific gravity to the new type of fluid that's been in use since 2011) still refuses to leave the system.

Draining the coupling fluid out of the gaps in the triplet
Here's the replacement fluid slowly going in...  As the gaps fill up, most of the evidence of the old fluid disappears again.

Lens fluid transfusion in progress
The finished product looks Really Fantastic!

Sparkly triplet with surfaces cleaned & lens fluid replaced
The throughput measurements were repeated after the fluid change & showed an additional 2-3% increase blueward of 3700 A.

Before cleaning (red), after cleaning (blue) & after fluid change (green) throughput measurements for the main group
Having completed the post-fluid-change measurements, the main group was re-installed in the instrument.

Putting the main group back in this evening
The main group & field lens assemblies go into the spectrograph from below.

Jono making sure everything's lined up properly
The lens barrel is supported on a wooden frame that's carefully positioned & then slowly jacked up on a lifting platform.

Almost there - soon to be bolted into position
With that, it was time to get started on the field lens group.  A relatively simple assembly consisting of a fused quartz bottom lens (L1) & a calcium fluoride one above it (L2), with a small fluid-filled gap in between.

The calcium fluoride (L2) lens in the RSS field lens assembly
The dirty assembly was rigged up to the throughput testing system, which consists of a broadband light source, an iris to control the beam size, various combinations of lenses (to re-image the source for each group to be tested), an integrating sphere to even out the light distribution & an extremely handy little spectrograph that Ted donated from his old lab.  The optic under test is mounted on a pneumatically actuated stage that places the lenses in & out of the beam to produce the differential measurement (a huge upgrade over the old system that relied on a person sitting in the dark actuating the stage!).  The even more significant improvement is the fact that the throughput measurements over a broad wavelength range are now in the form of spectra obtained at various grating angles, rather than photometric data taken with a set of manually deployed filters.  Janus has also written a data pipeline to painlessly calibrate the spectra & display the throughput results - brilliant! 

The new throughput testing system, with the RSS field lens group installed
Here's the profile view of the field lens group in the beam.

A bit of a tight fit
This is almost the photon's view of the setup...  The light source is in the lower right corner of the image & the lab spectrograph is visible in the background on the right.  A set of "before" measurements was taken this evening so that the cleaning can be done first thing in the morning...

Field lens group in beam in the throughput testing system
While this setting up was going on in the clean room, most of the team was still busy with various tasks in the spectrometer room - only about 11 hours after they showed up here this morning!

Tell us a story Uncle Paul!
The notoriously tricky slit-mask mechanism was putting Thabelo & co. through their paces...

Feeding the monster
2.5 weeks into the shut-down, things are still on schedule & going well - great work by all concerned!

No comments:

Post a Comment